Mattie drove 150 miles in the same amount of time that it took a turboprpellar plane to travel 600 miles. The spped of the plane was 150 mph faster than the speed of the car mattie drove. Find the speed of the plane.
This came from a take home exam from my Math Structures 1 teacher

Answer by checkley77(12569)

150/X=600/(X+150)
150(X+150)=600X
150X+22,500=600X
150X-600X=-22,500
-450X=-22,500
X=-22,500/-450
X=50 MPH IS MATTIE'S SPEED.
50=150=200 MPH IS THE PLANE'S SPEED.